# Dimension-independent convergence rate of propagation of chaos and numerical analysis for McKean-Vlasov stochastic differential equations with coefficients nonlinearly dependent on measure

### v1

@@ 4 Numerical simulation
Figure 3: $L^{4}$ convergence rate of the PoC.
(a) d = 10
(b) d = 12
(c) d = 14
Figure 4: $L^{6}$ convergence rate of the PoC.
Appendix A Proof of Theorem 2.1
First, let us introduce some definitions for simplicity and give some lemmas.
For $x\in\mathbb{R}^{d}$, define
$\hat{\kappa}(x):=\int_{\mathbb{R}^{d}}\kappa(x,y)\mu^{x}({\rm d}y),\quad\hat{\zeta}(x):=\int_{\mathbb{R}^{d}}\zeta(x,z)\mu^{x}({\rm d}z),$
then (5) equals to
Correspondingly, for $i=1,2,\dots,N$ and $x^{i}\in\mathbb{R}^{d}$, define
then (7) equals to
Lemma A.1.
Let $p>2$, for any $t\in[0,T]$, it holds under Assumption 2.3 that
$\mathbb{E}\left|\hat{\kappa}(x_{t}^{i})-\hat{\kappa}^{N}(x_{t}^{i})\right|^{p}\vee\mathbb{E}\left|\hat{\zeta}(x_{t}^{i})-\hat{\zeta}^{N}(x_{t}^{i})\right|^{p}\leq CN^{-p/2}.$
Proof.
According to inequality (4), and the Rosenthal inequality (see [36]), we can deduce that
$\leq$ $CN^{-p}\left(\mathbb{E}\left|\int_{\mathbb{R}^{d}}\left(\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{i})\right)\mu_{t}^{x^{i}}({\rm d}y)\right|^{p}+\mathbb{E}\left|\sum_{j\neq i}\int_{\mathbb{R}^{d}}\left(\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{j})\right)\mu_{t}^{x^{i}}({\rm d}y)\right|^{p}\right)$
$+CN^{-p}\left(\sum_{j\neq i}\mathbb{E}\left|\int_{\mathbb{R}^{d}}\left(\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{j})\right)\mu_{t}^{x^{i}}({\rm d}y)\right|^{2}\right)^{p/2}.$
Then, applying the Hölder inequality, Assumption 2.3 and Lemma 2.1, yields that
$\mathbb{E}|\hat{\kappa}(x_{t}^{i})-\hat{\kappa}^{N}(x_{t}^{i})|^{p}$
$\leq$ $CN^{-p}\mathbb{E}\int_{\mathbb{R}^{d}}\left|\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{i})\right|^{p}\mu_{t}^{x^{i}}({\rm d}y)+CN^{-p}\sum_{j\neq i}\mathbb{E}\int_{\mathbb{R}^{d}}\left|\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{j})\right|^{p}\mu_{t}^{x^{i}}({\rm d}y)$
$+CN^{-p}(N-1)^{p/2-1}\sum_{j\neq i}\mathbb{E}\int_{\mathbb{R}^{d}}\left|\kappa(x_{t}^{i},y)-\kappa(x_{t}^{i},x_{t}^{j})\right|^{p}\mu_{t}^{x^{i}}({\rm d}y)$
$\leq$ $CN^{-p}\mathbb{E}\int_{\mathbb{R}^{d}}|y-x_{t}^{i}|^{p}\mu_{t}^{x^{i}}({\rm d}y)+C(N^{-p}+N^{-p/2-1})\sum_{j\neq i}\mathbb{E}\int_{\mathbb{R}^{d}}|y-x_{t}^{j}|^{p}\mu_{t}^{x^{i}}({\rm d}y)$
$\leq$ $CN^{-p}\mathbb{E}\left(\mathbb{E}|x_{t}^{i}|^{p}+|x_{t}^{i}|^{p}\right)+C(N^{-p}+N^{-p/2-1})\sum_{j\neq i}\mathbb{E}\left(\mathbb{E}|x_{t}^{i}|^{p}+|x_{t}^{j}|^{p}\right)$
$\leq$ $CN^{-p/2}(N^{-p/2}+N^{-p/2+1}+1)$
$\leq$ $CN^{-p/2}.$ (17)
Analogously we can also get
$\mathbb{E}\left|\hat{\zeta}(x_{t}^{i})-\hat{\zeta}^{N}(x_{t}^{i})\right|^{p}\leq CN^{-p/2}$
from Assumption 2.3. ∎
Lemma A.2.
Let $p>2$, for any $t\in[0,T]$, it holds under Assumptions 2.1, 2.3 and 2.4 that
$\mathbb{E}|X_{t}^{i,N,\delta}-X_{\delta_{t}}^{i,N,\delta}|^{p}\leq C(t-\delta_{t})^{p/2}.$
Proof.
$X_{t}^{i,N,\delta}=$ $X_{\delta_{t}}^{i,N,\delta}+\int_{\delta_{t}}^{t}\tilde{a}(X_{\delta_{s}}^{i,N,\delta}){\rm d}s+\int_{\delta_{t}}^{t}A\left(\int_{\mathbb{R}^{d}}\kappa(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right){\rm d}s$
$+\int_{\delta_{t}}^{t}B\left(\int_{\mathbb{R}^{d}}\zeta(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right){\rm d}W_{s}^{i}.$
On the other hand, for $s\in[\delta_{t},t]$, since
and
we can get that $\delta_{s}=\delta_{t}$. Hence, using (4), the B-D-G inequality, and Remark 2.1, we have
$\mathbb{E}|X_{t}^{i,N,\delta}-X_{\delta_{t}}^{i,N,\delta}|^{p}$
$=$ $\mathbb{E}\left|\int_{\delta_{t}}^{t}\tilde{a}(X_{\delta_{s}}^{i,N,\delta}){\rm d}s+\int_{\delta_{t}}^{t}A\left(\int_{\mathbb{R}^{d}}\kappa(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right){\rm d}s+\int_{\delta_{t}}^{t}B\left(\int_{\mathbb{R}^{d}}\zeta(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right){\rm d}W_{s}^{i}\right|^{p}$
$\leq$ $C(t-\delta_{t})^{p}\mathbb{E}\left|\tilde{a}(X_{\delta_{t}}^{i,N,\delta})\right|^{p}+C(t-\delta_{t})^{p}\mathbb{E}\left|A\left(\int_{\mathbb{R}^{d}}\kappa(X_{\delta_{t}}^{i,N,\delta},y)\mu_{\delta_{t}}^{X,N,\delta}({\rm d}y)\right)\right|^{p}$
$+C(t-\delta_{t})^{p/2-1}\mathbb{E}\int_{\delta_{t}}^{t}\left|B\left(\int_{\mathbb{R}^{d}}\zeta(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right)\right|^{p}{\rm d}s$
$\leq$ $C(t-\delta_{t})^{p}\mathbb{E}\left|\tilde{a}(X_{\delta_{t}}^{i,N,\delta})\right|^{p}+C(t-\delta_{t})^{p}+C(t-\delta_{t})^{p}\mathbb{E}\left|\int_{\mathbb{R}^{d}}\kappa(X_{\delta_{t}}^{i,N,\delta},y)\mu_{\delta_{t}}^{X,N,\delta}({\rm d}y)\right|^{p}$
$+C(t-\delta_{t})^{p/2}\mathbb{E}\left|\int_{\mathbb{R}^{d}}\zeta(X_{\delta_{s}}^{i,N,\delta},y)\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)\right|^{p}$
$\leq$ $C(t-\delta_{t})^{p}\mathbb{E}\left|\tilde{a}(X_{\delta_{t}}^{i,N,\delta})\right|^{p}+C(t-\delta_{t})^{p}+C(t-\delta_{t})^{p}\mathbb{E}\int_{\mathbb{R}^{d}}\left|\kappa(X_{\delta_{t}}^{i,N,\delta},y)\right|^{p}\mu_{\delta_{t}}^{X,N,\delta}({\rm d}y)$
$+C(t-\delta_{t})^{p/2}\mathbb{E}\int_{\mathbb{R}^{d}}\left|\zeta(X_{\delta_{s}}^{i,N,\delta},y)\right|^{p}\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)$
$\leq$ $C(t-\delta_{t})^{p}\mathbb{E}\left|\tilde{a}(X_{\delta_{t}}^{i,N,\delta})\right|^{p}+C(t-\delta_{t})^{p}+C(t-\delta_{t})^{p/2}\mathbb{E}\int_{\mathbb{R}^{d}}(|X_{\delta_{t}}^{i,N,\delta}|^{p}+|y|^{p})\mu_{\delta_{s}}^{X,N,\delta}({\rm d}y)$
$\leq$ $C(t-\delta_{t})^{p/2}\left((t-\delta_{t})^{p/2}\mathbb{E}\left|\tilde{a}(X_{\delta_{t}}^{i,N,\delta})\right|^{p}+(t-\delta_{t})^{p/2}+\mathbb{E}|X_{\delta_{t}}^{i,N,\delta}|^{p}\right).$
Since $t-\delta_{t}\leq t\leq T$, applying Assumption 2.4, it holds that
$\mathbb{E}|X_{t}^{i,N,\delta}-X_{\delta_{t}}^{i,N,\delta}|^{p}\leq C(t-\delta_{t})^{p/2}.$
Proof of Theorem 2.1.
For any $i=1,2,\dots,N$, according to (15) and (16), we have
For any $i\in\{1,2,\dots,N\}$, applying the Itô formula, the Cauchy-Schwarz inequality, and Assumption 2.3, one has
$|e_{t}^{i,N}|^{p}\leq$ $p\int_{0}^{t}|e_{s}^{i,N}|^{p-2}\left\langle e_{s}^{i,N},a(x_{s}^{i})-\tilde{a}(X_{\delta_{s}}^{i,N,\delta})\right\rangle{\rm d}s+p\lambda_{2}\int_{0}^{t}|e_{s}^{i,N}|^{p-1}\left|\hat{\kappa}(x_{s}^{i})-\hat{\kappa}^{N}(X_{\delta_{s}}^{i,N,\delta})\right|{\rm d}s$
$+\frac{p(p-1)\lambda_{3}^{2}}{2}\int_{0}^{t}|e_{s}^{i,N}|^{p-2}\left|\hat{\zeta}(x_{s}^{i})-\hat{\zeta}^{N}(X_{\delta_{s}}^{i,N,\delta})\right|^{2}{\rm d}s$
$+p\lambda_{3}\int_{0}^{t}|e_{s}^{i,N}|^{p-1}\left|\hat{\zeta}(x_{s}^{i})-\hat{\zeta}^{N}(X_{\delta_{s}}^{i,N,\delta})\right|{\rm d}W_{s}^{i}$
=:absent:=:= : $J_{1}+J_{2}+J_{3}+J_{4}.$ (18)
$\leq$ $C\int_{0}^{t}|e_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}(1+|X_{s}^{i,N,\delta}|^{q}+|X_{\delta_{s}}^{i,N,\delta}|^{q})^{p}|X_{s}^{i,N,\delta}-X_{\delta_{s}}^{i,N,\delta}|^{p}{\rm d}s$
Similarly, we can also be deduced from the Young inequality and Assumption 2.1 that
$J_{2}\leq$ $p\lambda_{2}\int_{0}^{t}|e_{s}^{i,N}|^{p-1}\left(|\hat{\kappa}(x_{s}^{i})-\hat{\kappa}^{N}(x_{s}^{i})|+|\hat{\kappa}^{N}(x_{s}^{i})-\hat{\kappa}^{N}(X_{s}^{i,N,\delta})|+|\hat{\kappa}^{N}(X_{s}^{i,N,\delta})-\hat{\kappa}^{N}(X_{\delta_{s}}^{i,N,\delta})|\right){\rm d}s$
$\leq$ $(p-1)\lambda_{2}\int_{0}^{t}|e_{s}^{i,N}|^{p}{\rm d}s+\lambda_{2}\int_{0}^{t}|\hat{\kappa}(x_{s}^{i})-\hat{\kappa}^{N}(x_{s}^{i})|^{p}{\rm d}s+\frac{\lambda_{2}}{N}\sum_{j=1}^{N}\int_{0}^{t}\left(|e_{s}^{i,N}|^{p}+|e_{s}^{j,N}|^{p}\right){\rm d}s$
$+\frac{\lambda_{2}}{N}\sum_{j=1}^{N}\int_{0}^{t}\left(|X_{s}^{j,N,\delta}-X_{\delta_{s}}^{i,N,\delta}|^{p}+|X_{s}^{j,N,\delta}-X_{\delta_{s}}^{j,N,\delta}|^{p}\right){\rm d}s,$ (20)
and
$+\frac{C}{N}\sum_{j=1}^{N}\int_{0}^{t}\left(|X_{s}^{j,N,\delta}-X_{\delta_{s}}^{i,N,\delta}|^{p}+|X_{s}^{j,N,\delta}-X_{\delta_{s}}^{j,N,\delta}|^{p}\right){\rm d}s.$ (21)
Substituting (A)-(A) into (A), it gives that
$|e_{t}^{i,N}|^{p}\leq$ $C\int_{0}^{t}|e_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}(1+|X_{s}^{i,N,\delta}|^{q}+|X_{\delta_{s}}^{i,N,\delta}|^{q})^{p}|X_{s}^{i,N,\delta}-X_{\delta_{s}}^{i,N,\delta}|^{p}{\rm d}s$
$\mathbb{E}\left(\sup_{0\leq t\leq T_{1}}|e_{t}^{i,N}|^{p}\right)\leq$ $C\int_{0}^{T_{1}}\mathbb{E}|e_{t}^{i,N}|^{p}{\rm d}t+C\int_{0}^{T_{1}}\left(\mathbb{E}(1+|X_{t}^{i,N,\delta}|^{q}+|X_{\delta_{t}}^{i,N,\delta}|^{q})^{2p}\mathbb{E}|X_{t}^{i,N,\delta}-X_{\delta_{t}}^{i,N,\delta}|^{2p}\right)^{1/2}{\rm d}t$
$+C\mathbb{E}\sup_{0\leq t\leq T_{1}}\left(\int_{0}^{t}|e_{s}^{i,N}|^{p-1}\left|\hat{\zeta}(x_{s}^{i})-\hat{\zeta}^{N}(X_{s}^{i,N,\delta})\right|{\rm d}W_{s}^{i}\right).$ (22)
Using the B-D-G inequality, similar to (A), it holds that
$\mathbb{E}\sup_{0\leq t\leq T_{1}}\left(\int_{0}^{t}|e_{s}^{i,N}|^{p-1}\left|\hat{\zeta}(x_{s}^{i})-\hat{\zeta}^{N}(X_{\delta_{s}}^{i,N,\delta})\right|{\rm d}W_{s}^{i}\right)\leq C\mathbb{E}\left(\int_{0}^{T_{1}}|e_{t}^{i,N}|^{2p-2}|\hat{\zeta}(x_{t}^{i})-\hat{\zeta}^{N}(X_{\delta_{t}}^{i,N,\delta})|^{2}{\rm d}t\right)^{1/2}$
Thus, combining (A) and (A), one has
$\mathbb{E}\left(\sup_{0\leq t\leq T_{1}}|e_{t}^{i,N}|^{p}\right)\leq$ $C\int_{0}^{T_{1}}\mathbb{E}|e_{t}^{i,N}|^{p}{\rm d}t+C\int_{0}^{T_{1}}\left(\mathbb{E}(1+|X_{t}^{i,N,\delta}|^{q}+|X_{\delta_{t}}^{i,N,\delta}|^{q})^{2p}\mathbb{E}|X_{t}^{i,N,\delta}-X_{\delta_{t}}^{i,N,\delta}|^{2p}\right)^{1/2}{\rm d}t$
Substituting Lemmas A.1 and A.2 above into (A), using Assumption 2.4 again, it gives
$\mathbb{E}\left(\sup_{0\leq t\leq T_{1}}|e_{t}^{i,N}|^{p}\right)\leq$ $C\int_{0}^{T_{1}}\mathbb{E}|e_{t}^{i,N}|^{p}{\rm d}t+C\int_{0}^{T_{1}}(t-\delta_{t})^{p/2}\left(1+\left\{\mathbb{E}\left(\sup_{0\leq s\leq t}|X_{s}^{i,N,\delta}|^{2pq}\right)\right\}^{1/2}\right){\rm d}t$
$+CN^{-p/2}+\int_{0}^{T_{1}}\mathbb{E}|a(X_{\delta_{t}}^{i,N,\delta})-\tilde{a}(X_{\delta_{t}}^{i,N,\delta})|^{p}{\rm d}t$
$\leq$ $C\int_{0}^{T_{1}}\mathbb{E}\left(\sup_{0\leq s\leq t}|e_{s}^{i,N}|^{p}\right){\rm d}t+CN^{-p/2}+C\int_{0}^{T_{1}}(t-\delta_{t})^{p/2}{\rm d}t$
Therefore, it follows from the Gronwall inequality that
$\mathbb{E}\left(\sup_{0\leq t\leq T_{1}}|e_{t}^{i,N}|^{p}\right)\leq$ $CN^{-p/2}+C\sup_{0\leq t\leq T_{1}}(t-\delta_{t})^{p/2}+C\mathbb{E}\left(\sup_{0\leq t\leq T_{1}}|a(X_{t}^{i,N,\delta})-\tilde{a}(X_{t}^{i,N,\delta})|^{p}\right).$
Then the assertion follows. ∎

### v2

@@ 6 Numerical simulation
Figure 3: $L^{4}$ convergence rate of PoC for Example 2.
(a) d = 2
(b) d = 4
(c) d = 6
Figure 4: $L^{2}$ convergence rate of PoC for Example 3.
(a) d = 6
(b) d = 8
(c) d = 10
Figure 5: $L^{4}$ convergence rate of PoC for Example 3.
Example 3.
In this example, we test the rate of PoC for MV-SDEs with the more general nonlinear distribution-dependence coefficients. Consider the $d$-dim nonlinear MV-SDE (4) with coefficients $a(x)=x-x^{5}$, $A(x,y)=\frac{1}{1+{\rm e}^{-x-y}}$, $B(x,y)=\sqrt{x^{2}+y^{2}}$, $\kappa_{1}(x,y)=\arctan(x+y)$, $\kappa_{2}(x,y)=\arctan(x-y)$,
$\zeta_{1}(x,y)=$ $\left(\begin{array}[]{cccc}|x_{1}+y_{1}|&x_{2}&\cdots&x_{d}\\ x_{1}&|x_{2}+y_{2}|&\cdots&x_{d}\\ \vdots&\vdots&&\vdots\\ x_{1}&x_{2}&\cdots&|x_{d}+y_{d}|\end{array}\right),\quad\zeta_{2}(x,y)=\left(\begin{array}[]{cccc}|x_{1}-y_{1}|&x_{2}&\cdots&x_{d}\\ x_{1}&|x_{2}-y_{2}|&\cdots&x_{d}\\ \vdots&\vdots&&\vdots\\ x_{1}&x_{2}&\cdots&|x_{d}-y_{d}|\end{array}\right)$
let $x_{0}\sim N(0,1)$, the the tamed EM scheme is given by
$X_{t_{n+1}}^{i,N}=$ $X_{t_{n}}^{i,N}+\frac{X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{5}}{1+\sqrt{\Delta}|X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{5}|}\Delta+\left(1+\exp\left\{-\frac{1}{N}\sum_{j=1}^{N}\left(\kappa_{1}(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})-\kappa_{2}(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})\right)\right\}\right)^{-1}\Delta$
$+\sqrt{\left(\frac{1}{N}\sum_{j=1}^{N}\zeta_{1}(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})\right)^{2}+\left(\frac{1}{N}\sum_{j=1}^{N}\zeta_{2}(X_{t_{n}}^{i,N},X_{t_{n}}^{j,N})\right)^{2}}\Delta W_{n}^{i}.$
The orders of convergence with respect to the number of particles $N$ are shown in Figures 4 and 5. The proxy solution takes $N=2^{11}$. The numerical solutions take $\bar{N}=2^{7},2^{8},2^{9},2^{10}$.
Example 4.
Finally, we test the PoC rate for MV-SDEs with coefficients contain the 2nd order interaction terms. Constrained by the computational capability of our computers, here we only give numerical verification that the rate of convergence can reach $1/2$ for the following 1-dimensional equation. Let $a(x)=x-x^{5}$, $A(x)=\tanh(x)$, $B(x)=\frac{1}{1+{\rm e}^{-x}}$, $\kappa(x,y,z)=|x+y+z|$, $\zeta(x,y,z)=\frac{x+y}{\sqrt{1+x^{2}+y^{2}+z^{2}}}$. let $x_{0}\sim N(0,1)$, the tamed EM scheme is given by
$X_{t_{n+1}}^{i,N}=$ $X_{t_{n}}^{i,N}+\frac{X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{5}}{1+\sqrt{\Delta}|X_{t_{n}}^{i,N}-(X_{t_{n}}^{i,N})^{5}|}\Delta+\tanh\left(\frac{1}{N^{2}}\sum_{j,r=1}^{N}|X_{t_{n}}^{i,N}+X_{t_{n}}^{j,N}+X_{t_{n}}^{r,N}|\right)\Delta$
$+\left(1+\exp\left\{-\frac{1}{N^{2}}\sum_{j,r=1}^{N}\frac{X_{t_{n}}^{i,N}+X_{t_{n}}^{j,N}}{\sqrt{1+(X_{t_{n}}^{i,N})^{2}+(X_{t_{n}}^{j,N})^{2}+(X_{t_{n}}^{r,N})^{2}}}\right\}\right)^{-1}\Delta W_{n}^{i}.$
The orders of convergence with respect to the number of particles $N$ are shown in Figure 6. The proxy solution takes $N=2^{8}$. The numerical solutions take $\bar{N}=2^{4},2^{5},2^{6},2^{7}$.
Figure 6: $L^{p}$ convergence rate of PoC for Example 4.
7 Discussion
The main contribution of this paper is that we theoretically prove that the PoC convergence in the sense of $L^{p},p\geq 2$ is of order $\mathcal{O}(N^{-1/2})$, which is dimension-independent, for several class of McKean-Vlasov stochastic differential equations (MV-SDEs) with coefficients non-linear dependence on the measure component. Both theoretical and numerical results show that this rate of convergence is across dimensions. However, the equations considered in this work, while nonlinear in their dependence on the measure, still necessitate satisfying the Lipschitz conditions. When the dependence of the coefficients on the measure is highly nonlinear, the relevant research still deserves further exploration. In addition, an off-topic note, in view of the computational time and cost, the particle approximation does not seem to be a good choice when the equations contain higher-order interaction terms, and it would be more important to explore the numerical schemes with lower computational cost.